Winter 2019 PHYS 232-02 Problem 22.14 v Part A M Review What is the magnitude of the electric force on each sphere? Express your answer to two significant figures and include the appropriate units Two small plastic spheres each have a mass of 1.3 g and a charge of 60.0 nC. They are placed 2.4 cm apart (center to center) 0図 F- 1 Value Units Submit Request Answer Part B By what factor is the electric force on a sphere larger than its weight? Express your answer using two significant figures. 画? ng Submit Provide Feedback 0 Type here to search
Answer #1

Part A.

Electrostatic force is given by:

F = k*q1*q2/r^2

k = electrostatic constant = 9*10^9

q1 = q2 = charge on each sphere = -60 nC = -60*10^-9 C

r = distance between both spheres = 2.4 cm = 0.024 m

So,

|F| = k*|q1|*|q2|/r^2

|F| = 9*10^9*60*10^-9*60*10^-9/0.024^2

|F| = 0.05625 N

In two significant figures

|F| = 5.6*10-2 N = 5.6*10^-2 N

Part B.

Now we need to find F/mg = ?

m = mass of sphere =1.3 gm = 1.3*10^-3 kg

g = gravitational acceleration = 9.81 m/sec^2

So,

F/mg = 0.05625/(1.3*10^-3*9.81)

F/mg = 4.4

Please Upvote

Comment below if you have any query.